OpenStudy (anonymous):

two times a number plus three equals one half of the number plus 12.what is the number? a.3.6 b.6 c.8 i.10 please work out problem

OpenStudy (anonymous):

Let's call the number we're looking for x. Then we would know that the following must hold for x:\[2x + 3 = \frac{x}{2} + 12\]Now you just have to solve this for x. Can you do that on your own?

OpenStudy (anonymous):

can you just give me the problem worked out with the answer cause im in a test now

OpenStudy (anonymous):

\[2x - \frac{x}{2} = 12 - 3\]\[\frac{3}{2}x = 9\]\[3x = 18\]\[x = 6\]Lol. Better try learning this stuff next time. It's not really hard and you're probably going to need it anyway later on. Good luck for your test

OpenStudy (anonymous):

6=-3(x+4)

OpenStudy (anonymous):

6 = -3x -12 3x = -18 x = 6

OpenStudy (anonymous):

@susanc27 that's not right. if 3x = - 18, then x = -6.

OpenStudy (anonymous):

Right. Typo. Thanks for the correction.